The density of iron is 7860 kg/m^3 what is the mass of an iron sphere whose diameter is 0.50 m


1
Expert's answer
2021-07-25T09:17:16-0400
"m=\rho V=\dfrac{4}{3}\pi r^3\rho,""m=\dfrac{4}{3}\pi\cdot(0.25\ m)^3\cdot7860\ \dfrac{kg}{m^3}=514\ kg."